OpenStudy (anonymous):

6x - 8 + 2x - 5 = 7 - 2x

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

find X please

OpenStudy (anonymous):

6x - 8 + 8 + 2x - 5 = 7 - 2x +8 6x + 2x - 5 = 7 - 2x +8 6x + 2x - 5 + 5= 7 - 2x +8 + 5 6x + 2x = 7 - 2x +8 + 5 6x + 2x + 2x= 7 - 2x +8 + 5 + 2x 6x + 2x + 2x= 7 +8 + 5